# Hierarchy of timescales in a disordered spin-$1/2$ XX ladder

### v2

@@ Acknowledgements.
The data and code for this paper are available on Zenodo [141].
@@ Appendix A Disorder-induced delocalized-localized regime crossover diagram
Figure 9: Disorder-induced delocalization-localization crossover diagrams of a spin-$1/2$ XX ladder as function of the disorder strength $W/J_{\parallel}$ and the coupling ratio $J_{\perp}/J_{\parallel}$ for $L=9$ for an energy density in the center of the spectrum. The diagrams are based on two different measures: (a) the state-to-state and sample-to-sample average of the adjacent gap ratio $\overline{r}$, and (b) the state-to-state and sample-to-sample average of the standard deviation of the entanglement entropy, $\sigma_{\overline{S}_{\mathrm{vNE}}}$. Each colored rectangle represents an average over two steps: first, a subset of eigenstates around the center of the energy spectrum, and then $M=90$ disorder realizations. In (b), the red dots indicate the maximum value of $\sigma_{\overline{S}_{\mathrm{vNE}}}{(W/J_{\parallel})}$ at a fixed value of $J_{\perp}/J_{\parallel}$.
@@ A.1 Adjacent gap ratio
The adjacent gap ratio allows us to check whether the spectral level statistics of the system follow the predictions from the RMT. These predictions purely originate from the adjacent energy gap $\omega_{E,\alpha}{(\boldsymbol{\mathcal{W}})}$, defined as
$\omega_{E,\alpha}{(\boldsymbol{\mathcal{W}})}\coloneqq E_{\alpha}{(\boldsymbol{\mathcal{W}})}-E_{\alpha-1}{(\boldsymbol{\mathcal{W}})}\,$, (38)
where $\alpha$ is the index for the larger eigenvalue of the adjacent energy pair $\{E_{\alpha-1},E_{\alpha}\}$. We then introduce the dimensionless adjacent gap ratio for a single disorder realization, $r_{\alpha}{(\boldsymbol{\mathcal{W}})}$ as follows:
$r_{\alpha}{(\boldsymbol{\mathcal{W}})}=\frac{\min\{\omega_{E,\alpha}{(\boldsymbol{\mathcal{W}})},\omega_{E,\alpha+1}{(\boldsymbol{\mathcal{W}})}\}}{\max\{\omega_{E,\alpha}{(\boldsymbol{\mathcal{W}})},\omega_{E,\alpha+1}{(\boldsymbol{\mathcal{W}})}\}}\,.$ (39)
The motivation for utilizing the ratios of energy gaps rather than the energy gaps themselves originates from the need of unfolding when using the latter.
Due to the potential edge effects in the energy spectrum, which arise from the very low density of states at extreme energy eigenvalues, we only use some portion of the energy eigenvalues from the spectral center for the adjacent gap ratio. We introduce the energy density $\tilde{E}_{\alpha}{(\boldsymbol{\mathcal{W}})}$, given by
$\tilde{E}_{\alpha}{(\boldsymbol{\mathcal{W}})}=\frac{E_{\alpha}{(\boldsymbol{\mathcal{W}})}-E_{\mathrm{min}}{(\boldsymbol{\mathcal{W}})}}{E_{\mathrm{max}}{(\boldsymbol{\mathcal{W}})}-E_{\mathrm{min}}{(\boldsymbol{\mathcal{W}})}}\,$, (40)
which reduces the total energy interval to $[0,1]$. For the ladder model from Eq. (1), we only consider $10\%$ closest eigenstates to the center of the energy spectrum in the $S^{z}=0$ subspace.
Based on Eqs. (38) and (39), the state-to-state and sample-to-sample average of adjacent gap ratio, $\overline{r}$, amounts to
$\overline{r}=\mathbb{E}_{\boldsymbol{\mathcal{W}}}\quantity[\mathbb{E}_{\mathrm{c}}\quantity[r_{\alpha}{(\boldsymbol{\mathcal{W}})}]]\,$, (41)
where we first carry out the central-portion and disorder averages in succession.
@@ A.2 Standard deviation of the entanglement entropy
In the delocalized regime and the localized regime, $\sigma_{\overline{S}_{\mathrm{vNE}}}$ converges towards zero as the system size increases. At the onset of crossover regime, $\sigma_{\overline{S}_{\mathrm{vNE}}}$ exhibits a peak whose amplitude increases and whose width decreases with increasing system size.
In Fig. 9(b), we show the state-to-state and sample-to-sample average of the standard devitation of the von Neumann entanglement entropy, $\sigma_{\overline{S}_{\mathrm{vNE}}}$, as a function of both the disorder strength $W/J_{\parallel}$ and the coupling ratio $J_{\perp}/J_{\parallel}$. We observe the same behavior as for the gap ratio shown Fig. 9(a). For our study, we choose $W/J_{\parallel}<2$, which puts us safely away from the finite-size localized regime.
Appendix B Spectral form factor (SFF)
@@ B.1 Unfolding the energy spectrum
$G{\quantity(E;\boldsymbol{\mathcal{W}})}\coloneqq\sum_{\alpha=1}^{\mathcal{D}}\Theta{\quantity[E-E_{\alpha}{\quantity(\boldsymbol{\mathcal{W}})}]}\,$, (45)
where $\Theta{(\,\cdots)}$ denotes the Heaviside step function, and $\mathcal{D}$ is the Hilbert-space dimension. We then smoothen out this function based on the polynomial regression with the degree $n$, resulting in an approximating function $\hat{G}^{(n)}(E;\boldsymbol{\mathcal{W}})$. This function is used to define the $n$-th-degree unfolded energy eigenvalues $\varepsilon^{(n)}_{\alpha}{(\boldsymbol{\mathcal{W}})}$, given by
$\varepsilon^{(n)}_{\alpha}{(\boldsymbol{\mathcal{W}})}=\hat{G}^{(n)}{(E_{\alpha};\boldsymbol{\mathcal{W}})}\,.$ (46)
@@
$\Lambda^{(n)}{\quantity(\boldsymbol{\mathcal{W}})}=\frac{4}{\mathcal{D}}\sum_{\alpha=1}^{\mathcal{D}}\quantity[\frac{\varepsilon^{(n)}_{\alpha}{(\boldsymbol{\mathcal{W}})}-G{\quantity(E_{\alpha};\boldsymbol{\mathcal{W}})}}{\varepsilon^{(n)}_{\alpha}{(\boldsymbol{\mathcal{W}})}+G{\quantity(E_{\alpha};\boldsymbol{\mathcal{W}})}}]^{2}\,.$ (47)
This criterion function measures how the unfolded energies for a chosen polynomial degree deviate from the cumulative spectral function defined in Eq. (45). By searching in a certain interval for the polynomial degrees, the optimum polynomial degree for unfolding, $n_{0}{\quantity(\boldsymbol{\mathcal{W}})}$, can be found depending on the criterion function in Eq. (47). In this paper, we limit the polynomial degrees to be in the interval $[2,18]$, so that
$n_{0}{\quantity(\boldsymbol{\mathcal{W}})}\coloneqq\operatorname*{arg\,min}_{2\leq n\leq 18}{\Lambda^{(n)}{\quantity(\boldsymbol{\mathcal{W}})}}\,.$ (48)
Based on the optimal polynomial degree given in Eq. (48), the (optimal) unfolded energy eigenvalue for a disorder realization, denoted by $\varepsilon_{\alpha}{(\boldsymbol{\mathcal{W}})}$, can be expressed as
$\varepsilon_{\alpha}{(\boldsymbol{\mathcal{W}})}=\varepsilon^{(n_{0})}_{\alpha}{(\boldsymbol{\mathcal{W}})}\,.$ (49)
@@ B.2 Gaussian filtering
In addition to the spectral unfolding procedure covered in Appendix B.1, we also apply a Gaussian filter function to mitigate the edge effects arising in $\overline{K}{(\tau)}$, as defined in Eq. (5), as a result of the unfolding procedure itself. We define this filter function for an unfolded energy eigenvalue in a disorder realization, $\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha};\boldsymbol{\mathcal{W}})}$, as follows:
[…]
where $\overline{\varepsilon}{(\boldsymbol{\mathcal{W}})}$ and $\sigma_{\varepsilon}{(\boldsymbol{\mathcal{W}})}$ are the mean and the standard deviation of a given unfolded energy spectrum, $\eta$ is a factor that controls the width of the Gaussian filter, thereby tuning its filtering strength.
@@ B.3 Fixing the shift in the normalized Heisenberg time
To fix the shifting problem of the normalized Heisenberg time disccused in Subsec. III.1, we reweight each adjacent unfolded-energy gap by utilizing the filter function given in Appendix B.2, yielding the weighted-mean-level spacing
$\tilde{\delta\varepsilon}_{\mathrm{mean}}{(\boldsymbol{\mathcal{W}})}$ $\coloneqq$ $\tilde{N}_{\omega}^{-1}{(\boldsymbol{\mathcal{W}})}\sum_{\alpha}\omega_{\varepsilon,\alpha}{(\boldsymbol{\mathcal{W}})}$
$\times\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha};\boldsymbol{\mathcal{W}})}\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha-1};\boldsymbol{\mathcal{W}})}\,$,
where $\tilde{N}_{\omega}{(\boldsymbol{\mathcal{W}})}\coloneqq\sum_{\alpha}\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha};\boldsymbol{\mathcal{W}})}\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha-1};\boldsymbol{\mathcal{W}})}$ is the weighted number of adjacent energy gaps with respect to the filter $\rho_{\mathrm{f}}$. We then rescale the normalized times of each realization, denoted by $\tau{(\boldsymbol{\mathcal{W}})}$, with respect to $\tilde{\delta\varepsilon}_{\mathrm{mean}}{(\boldsymbol{\mathcal{W}})}$, i.e.,
$\tau{(\boldsymbol{\mathcal{W}})}\leftarrow\tilde{\delta\varepsilon}_{\mathrm{mean}}{(\boldsymbol{\mathcal{W}})}\times\tau{(\boldsymbol{\mathcal{W}})}\,.$ (52)
@@ B.4 Gaussian Orthogonal Ensemble
The analytical results for the spectral form factor (SFF) vary depending on the specific ensemble from the RMT being considered, but they can still be obtained analytically. For instance, in the case of a chaotic Hamiltonian following the time-reversal symmetry, the random matrix distributions correspond to the Gaussian Orthogonal Ensemble (GOE). The SFF for the GOE is given by [60]
@@ B.5 Heisenberg time
$\overline{\delta E}=\mathbb{E}_{\boldsymbol{\mathcal{W}}}\quantity[N_{\omega}^{-1}\sum_{\alpha}\omega_{E,\alpha}{(\boldsymbol{\mathcal{W}})}]\,$, (55)
@@ C.1 Initial current autocorrelation functions
With the local densities of transport quantities and local current operators defined as described in Subsec. III.2, we can proceed to establish the current autocorrelation function $C^{(\mu)}{(t;{\boldsymbol{\mathcal{W}},\beta})}$ within the framework of linear-response theory, as expressed in Eq. (24). In the limit of infinite temperature, where $\beta\rightarrow 0$, the current autocorrelation function simplifies to
@@
where $\mathcal{D}$ is the Hilbert-space dimension. This simplified expression provides a foundation for further analysis of the system’s transport properties in the high-temperature limit.
@@
$C^{\mathrm{(S)}}{(t=0;{\boldsymbol{\mathcal{W}}},\beta=0)}=\frac{J_{\parallel}^{2}}{8}\,$, (58)
@@
$C^{\mathrm{(E)}}{(t=0;{\boldsymbol{\mathcal{W}}},\beta=0)}=\frac{J_{\parallel}^{2}}{64}\quantity[2J_{\parallel}^{2}+J_{\perp}^{2}+\frac{2}{N}\sum_{\ell,i}\quantity(w_{\ell,i}+w_{\ell+1,i})^{2}]\,.$ (59)
@@ C.2 Static susceptibility per spin
$\overline{\chi_{\mu}}{(\beta=0)}=\mathbb{E}_{\boldsymbol{\mathcal{W}}}\quantity[\chi_{\mu}{(\boldsymbol{\mathcal{W}},\beta=0)}]\,$, (60)
where $\chi_{\mu}{(\boldsymbol{\mathcal{W}},\beta)}$ represents the static susceptibility per spin for a specific disorder realization at a given inverse temperature. This quantity is calculated using the following expression
$\chi_{\mu}{(\boldsymbol{\mathcal{W}},\beta)}=\frac{\expectationvalue{\quantity[Q^{(\mu)}{(\boldsymbol{\mathcal{W}})}]^{2}}_{\beta}-\expectationvalue{Q^{(\mu)}{(\boldsymbol{\mathcal{W}})}}_{\beta}^{2}}{N}\,.$ (61)
To obtain the disorder-averaged infinite-temperature static susceptibilities given in Eqs. (28) and (29), we first rewrite their conserved quantities as a linear combinations of Pauli strings for a single disorder realization, where every parameters of Hamiltonian are stored symbolically, and then trace them out to find their expectation values. For the disorder average, we symbolically integrate them over each $w_{\ell,i}$ with respect to their distributions, which eventually yields $\overline{\chi_{\mu}}{(\beta=0)}$.
C.3 Dynamical quantum typicality
The concept of dynamical quantum typicality (DQT) leverages the fact that, in a high-dimensional Hilbert space, a single pure state can replicate the characteristics of an entire statistical ensemble. This thus enables an approximate computation of the expectation value of current autocorrelation functions [129, 130, 131, 132, 133].
@@ C.3 Dynamical quantum typicality
To further simplify the time evolution in Eq. (62), we introduce two new states:
@@
These states represent the time-evolved thermal random states, with and without the application of the total current operator, respectively. Here, by switching from the Heisenberg picture to the Schrödinger picture, we basically express the time evolution in a more convenient form. Using Eqs. (63) and (64), we can approximate the current autocorrelation function $C^{(\mu)}{(t;{\boldsymbol{\mathcal{W}}},\beta)}$ as
@@
Figure 10: Time-dependent disorder-averaged spin-diffusion constants $\overline{D^{(\mathrm{S})}}{(t;\boldsymbol{\mathcal{W}})}$ of a spin-$1/2$ XX ladder as a function of time $t$ at disorder strength $W/J_{\parallel}=1.5$ and coupling ratio $J_{\perp}/J_{\parallel}=1.0$ for different numbers of rungs $L=4$ (blue), $L=5$ (yellow) and $L=6$ (green). Results obtained with exact diagonalization (ED) are shown as solid lines, while those from the dynamical quantum typicality (DQT) using the Lanczos algorithm are plotted as dashed lines. The parameters employed are identical to those listed in Fig. 6.
To assess the reliability of the diffusion constants obtained with the DQT, we compare them against exact‑diagonalization (ED) results in Fig. 10. As $L$ increases, the time-dependent disorder-averaged spin-diffusion constants computed via the DQT with the Lanczos algorithm exhibit a closer correspondence with the ED data. While for small $L\sim 4$, the data quality could be improved by sampling over more independent Haar-random states disorder realization, we find that going to larger $L$ the discrepancy with ED decreases quickly. [^4] We also verify that the utilization of same Haar-random state per disorder realization still gives rise to similar results. In our calculations for diffusion constants, we go up to $L=10$.

### v3

@@ Acknowledgements.
The data that support the findings of this article are partially openly available [140].
@@ Appendix A Disorder-induced delocalized-localized regime crossover diagram
Figure 9: Disorder-induced delocalization-localization crossover diagrams of a spin-$1/2$ XX ladder as function of the disorder strength $W/J_{\parallel}$ and the coupling ratio $J_{\perp}/J_{\parallel}$ for $L=9$ for a dimensionless normalized energy in the center of the spectrum. The diagrams are based on two different measures: (a) the state-to-state and sample-to-sample average of the adjacent gap ratio $\overline{r}$ and (b) the state-to-state and sample-to-sample average of the standard deviation of the entanglement entropy $\sigma_{\overline{S}_{\mathrm{vNE}}}$. Each colored rectangle represents an average over two steps: first, a subset of eigenstates around the center of the energy spectrum and then $M=90$ disorder realizations. In (b), the red dots indicate the maximum value of $\sigma_{\overline{S}_{\mathrm{vNE}}}{(W/J_{\parallel})}$ at a fixed value of $J_{\perp}/J_{\parallel}$.
@@ A.1 Adjacent gap ratio
The adjacent gap ratio allows us to check whether the spectral level statistics of the system follow the predictions from the RMT. These predictions purely originate from the adjacent energy gap $\omega_{\alpha}$, defined as
$\omega_{\alpha}\coloneqq E_{\alpha}-E_{\alpha-1}\,$, (37)
where $\alpha$ is the index for the larger eigenvalue of the adjacent energy pair $\{E_{\alpha-1},E_{\alpha}\}$. We then introduce the dimensionless adjacent gap ratio for a single disorder realization, $r_{\alpha}$ as follows
$r_{\alpha}=\frac{\min\{\omega_{\alpha},\omega_{\alpha+1}\}}{\max\{\omega_{\alpha},\omega_{\alpha+1}\}}\,.$ (38)
The motivation for utilizing the ratios of energy gaps rather than the energy gaps themselves results from the necessity to unfold the spectrum when using the latter.
Due to the potential edge effects in the energy spectrum, which arise from the very low density of states at extreme energy eigenvalues, we only use some portion of the energy eigenvalues from the spectral center for the adjacent gap ratio. We introduce the dimensionless normalized energy $\tilde{E}_{\alpha}$, given by
$\tilde{E}_{\alpha}=\frac{E_{\alpha}-E_{\mathrm{min}}}{E_{\mathrm{max}}-E_{\mathrm{min}}}\,$, (39)
which reduces the total energy interval to $[0,1]$. For the ladder model from Eq. (1), we only consider the $10\%$ closest eigenstates to the center of the energy spectrum in the $S^{z}=0$ subspace.
Based on Eqs. (37) and (38), the state-to-state and sample-to-sample average of adjacent gap ratio $\overline{r}$ amounts to
$\overline{r}=\mathbb{E}_{\boldsymbol{\mathcal{W}}}\quantity[\mathbb{E}_{\mathrm{c}}\quantity[r_{\alpha}]]\,$, (40)
where we first carry out the central-portion and disorder averages in succession denoted by $\mathbb{E}_{\mathrm{c}}$ and $\mathbb{E}_{\boldsymbol{\mathcal{W}}}$, respectively.
@@ A.2 Standard deviation of the entanglement entropy
In the delocalized regime and the localized regime, $\sigma_{\overline{S}_{\mathrm{vNE}}}$ converges toward zero as the system size increases. At the onset of crossover regime, $\sigma_{\overline{S}_{\mathrm{vNE}}}$ exhibits a peak whose amplitude increases and whose width decreases with increasing system size.
In Fig. 9(b), we show the state-to-state and sample-to-sample average of the standard devitation of the von Neumann entanglement entropy $\sigma_{\overline{S}_{\mathrm{vNE}}}$ as a function of both the disorder strength $W/J_{\parallel}$ and the coupling ratio $J_{\perp}/J_{\parallel}$. We observe the same behavior as for the gap ratio shown Fig. 9(a). For our study, we choose $W/J_{\parallel}<2$, which puts us safely away from the finite-size localized regime.
Appendix B SFF
@@ B.1 Unfolding the energy spectrum
$G{\quantity(E)}\coloneqq\sum_{\alpha=1}^{\mathcal{D}}\Theta{\quantity(E-E_{\alpha})}\,$, (44)
where $\Theta{(\,\cdots)}$ denotes the Heaviside step function, and $\mathcal{D}$ is the Hilbert-space dimension. We then smoothen out this function based on the polynomial regression with the degree $n$, resulting in an approximating function $\hat{G}^{(n)}(E)$. This function is used to define the $n$ th-degree unfolded energy eigenvalues $\varepsilon^{(n)}_{\alpha}$, given by
$\varepsilon^{(n)}_{\alpha}=\hat{G}^{(n)}{(E_{\alpha})}\,.$ (45)
@@
$\Lambda^{(n)}=\frac{4}{\mathcal{D}}\sum_{\alpha=1}^{\mathcal{D}}\quantity[\frac{\varepsilon^{(n)}_{\alpha}-G{\quantity(E_{\alpha})}}{\varepsilon^{(n)}_{\alpha}+G{\quantity(E_{\alpha})}}]^{2}\,.$ (46)
This criterion function measures how the unfolded energies for a chosen polynomial degree deviate from the cumulative spectral function defined in Eq. (44). By searching in a certain interval for the polynomial degrees, the optimum polynomial degree for unfolding $n_{0}{\quantity(\boldsymbol{\mathcal{W}})}$ can be found depending on the criterion function in Eq. (46). In this paper, we limit the polynomial degrees to be in the interval $[2,18]$, so that
$n_{0}\coloneqq\operatorname*{arg\,min}_{2\leq n\leq 18}{\Lambda^{(n)}}\,.$ (47)
Based on the optimal polynomial degree given in Eq. (47), the (optimal) unfolded energy eigenvalue for a disorder realization, denoted by $\varepsilon_{\alpha}$, can be expressed as
$\varepsilon_{\alpha}=\varepsilon^{(n_{0})}_{\alpha}\,$. (48)
@@ B.2 Gaussian filtering
In addition to the spectral unfolding procedure covered in Appendix B.1, we also apply a Gaussian filter function to mitigate the edge effects arising in $\overline{K}{(\tau)}$, as defined in Eq. (5), as a result of the unfolding procedure itself. We define this filter function for an unfolded energy eigenvalue in a disorder realization $\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha})}$ as follows:
$\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha})}=\exp[-\frac{1}{2}\quantity(\frac{\varepsilon_{\alpha}-\overline{\varepsilon}}{\eta\sigma_{\varepsilon}})^{2}]\,$, (49)
where $\overline{\varepsilon}$ and $\sigma_{\varepsilon}$ are the mean and the standard deviation of a given unfolded energy spectrum, and $\eta$ is a factor that controls the width of the Gaussian filter, thereby tuning its filtering strength.
@@ B.3 Fixing the shift in the normalized Heisenberg time
To fix the shifting problem of the normalized Heisenberg time discussed in Sec. III.1, we reweight each adjacent unfolded energy gap by utilizing the filter function given in Appendix B.2, yielding the weighted-mean-level spacing
$\overline{\tilde{\delta\varepsilon}}_{\mathrm{mean}}\coloneqq\mathbb{E}_{\boldsymbol{\mathcal{W}}}\quantity[\tilde{N}_{\omega}^{-1}\sum_{\alpha}\omega_{\alpha}\,\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha})}\,\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha-1})}]\,$, (50)
where $\tilde{N}_{\omega}\coloneqq\sum_{\alpha}\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha})}\,\rho_{\mathrm{f}}{\quantity(\varepsilon_{\alpha-1})}$ is the weighted number of adjacent energy gaps with respect to the filter $\rho_{\mathrm{f}}$. We then rescale the normalized times, denoted by $\tau$, with respect to $\overline{\tilde{\delta\varepsilon}}_{\mathrm{mean}}$, i.e.,
$\tau\leftarrow\overline{\tilde{\delta\varepsilon}}_{\mathrm{mean}}\times\tau\,.$ (51)
@@ B.4 Gaussian orthogonal ensemble
The analytical results for the SFF vary depending on the specific ensemble from the RMT being considered, but they can still be obtained analytically. For instance, in the case of a chaotic Hamiltonian following the time-reversal symmetry, the random matrix distributions correspond to the GOE. The SFF for the GOE is given by [59]
@@ B.5 Heisenberg time
$\overline{\delta E}=\mathbb{E}_{\boldsymbol{\mathcal{W}}}\quantity[N_{\omega}^{-1}\sum_{\alpha}\omega_{\alpha}]\,$, (54)
@@ C.1 Initial current autocorrelation functions
With the local densities of transport quantities and local current operators defined as described in Sec. III.2, we can proceed to establish the current autocorrelation function $C^{(\mu)}{(t;{\beta})}$ within the framework of linear-response theory, as expressed in Eq. (23). In the limit of infinite temperature, where $\beta\rightarrow 0$, the current autocorrelation function simplifies to
@@
where $\mathcal{D}$ is the Hilbert-space dimension and $N$ is the total number of spins. This simplified expression provides a foundation for further analysis of the transport properties of the system in the high-temperature limit.
@@
$C^{\mathrm{(S)}}{(t=0,\beta=0)}=\frac{J_{\parallel}^{2}}{8}\,$, (57)
@@
$C^{\mathrm{(E)}}{(t=0;\beta=0)}=\frac{J_{\parallel}^{2}}{64}\quantity[2J_{\parallel}^{2}+J_{\perp}^{2}+\frac{2}{N}\sum_{\ell,i}\quantity(w_{\ell,i}+w_{\ell+1,i})^{2}]\,.$ (58)
@@ C.2 Static susceptibility per spin
$\overline{\chi_{\mu}}{(\beta=0)}=\mathbb{E}_{\boldsymbol{\mathcal{W}}}\quantity[\chi_{\mu}{(\beta=0)}]\,$, (59)
where $\chi_{\mu}{(\beta)}$ represents the static susceptibility per spin for a specific disorder realization at a given inverse temperature. This quantity is calculated using the following expression
$\chi_{\mu}{(\beta)}=\frac{\expectationvalue{\quantity[Q^{(\mu)}]^{2}}_{\beta}-\expectationvalue{Q^{(\mu)}}_{\beta}^{2}}{N}\,.$ (60)
To obtain the disorder-averaged infinite-temperature static susceptibilities given in Eqs. (27) and (28), we first rewrite their conserved quantities as a linear combinations of Pauli strings for a single disorder realization, where every parameter of the Hamiltonian is stored symbolically, and then trace them out to find their expectation values. For the disorder average, we symbolically integrate them over each $w_{\ell,i}$ with respect to their distributions, which eventually yields $\overline{\chi_{\mu}}{(\beta=0)}$.
C.3 DQT
The concept of DQT leverages the fact that, in a high-dimensional Hilbert space, a single pure state can replicate the characteristics of an entire statistical ensemble. This thus enables an approximate computation of the expectation value of current autocorrelation functions [131, 132, 133, 134, 135].
@@ C.3 DQT
To further simplify the time evolution in Eq. (61), we introduce two new states
@@
These states represent the time-evolved thermal random states, with and without the application of the total current operator, respectively. Here, by switching from the Heisenberg picture to the Schrödinger picture, we basically express the time evolution in a more convenient form. Using Eqs. (62) and (63), we can approximate the current autocorrelation function $C^{(\mu)}{(t;\beta)}$ as
@@
Figure 10: Time-dependent disorder-averaged spin-diffusion constants $\overline{D^{(\mathrm{S})}}{(t)}$ of a spin-$1/2$ XX ladder as a function of time $t$ at disorder strength $W/J_{\parallel}=1.5$ and coupling ratio $J_{\perp}/J_{\parallel}=1.0$ for different numbers of rungs $L=4$ (blue), $L=5$ (yellow) and $L=6$ (green). Results obtained with ED are shown as solid lines, while those from the DQT using the Lanczos algorithm are plotted as dashed lines. The parameters employed are identical to those listed in Fig. 6.
To assess the reliability of the diffusion constants obtained with the DQT, we compare them against exact‑diagonalization (ED) results in Fig. 10. As $L$ increases, the time-dependent disorder-averaged spin-diffusion constants computed via the DQT with the Lanczos algorithm exhibit a closer correspondence with the ED data. While for small $L\sim 4$, the data quality could be improved by sampling over more independent Haar-random states disorder realization, we find that going to larger $L$ the discrepancy with ED decreases quickly. [^4] We also verify that the utilization of the same Haar-random state per disorder realization still gives rise to similar results. In our calculations for diffusion constants, we go up to $L=10$.